# “I think this is fair”: Uncovering the Complexities of Stakeholder Decision-Making in AI Fairness Assessment

### v1

@@ 4.2.2. Context Matters: Stakeholders Switch Metrics Across Features
Next, we analyzed each participant’s metric selections. We found that 20 participants often switched metrics for different features. For example, 10 participants used 3 different metrics, 8 used 2 metrics, and 2 used 4 metrics. An illustrative example is P7, who ranked and evaluated four distinct features, each aligned with a unique metric: (Consistency — ‘Credit History’); (Custom Fairness— ‘Purpose’); (Equal Opportunity — ‘Foreign Worker’); (Conditional Statistical Parity — ‘Gender’). We observed that participants seemed to first consider their preferred fairness category, i.e., individual or/and group fairness, and then decided on specific metrics within these categories. For example, P16 explained, “so to me the group fairness doesn’t take that much priority because the individual fairness is what matters” which then led to pairing different metrics within the individual fairness category with the selected features. Only six participants followed a “one-size-fits-all” metric selection approach, similar to the one currently adopted by AI experts, where they chose the same metric (or metric combinations under Custom Fairness) across all of the features they had selected.
@@ 4.2.3. Metrics for Features, and the Reasons Behind Them
We found that each of the top ranked features (‘Gender’, ‘Foreign Worker’, ‘Age’, ‘Telephone’, ‘Dependents’, and ‘Purpose’) were paired with multiple metrics: at least five different metrics in the case of ‘Telephone’, ‘Dependents’, and ‘Purpose’, to seven metrics for ‘Gender’. Since participants chose different metrics for different features, we investigated any patterns within feature-metric pairings, concentrating on the six most selected features and their most frequently selected metrics (Figure 4).
@@
Finally, we found that when predefined metrics did not meet participants’ perceptions, they were trying to define their own fairness metrics. Custom Fairness emerged as one of the most selected metrics for ‘Gender’ (7 participants), ‘Dependents’ (3 participants), and ‘Purpose’ (3 participants). Participants for the most part tried to combine different existing metrics to achieve a more comprehensive fairness assessment, for example by combining “Consistency and Counterfactual Fairness” for ‘Gender’ (2 participants) or “Equal Opportunity and Counterfactual Fairness” for ‘Gender’ (2 participants), ‘Dependents’ (2 participants) and ‘Purpose’ (2 participants). In constructing these custom metrics, such as “Consistency & Counterfactual Fairness”, participants wanted to meet both individual fairness metrics simultaneously. This, in their mind, would balance Comparison with similar individuals and Remove the influence of irrelevant features. Participants also mentioned the similarity between these two metrics. As P16 explained: “I think it’s interesting because I feel like the two things are connected so I felt like if counterfactual fairness talks about having the same result even when those specific features changed, and then consistency talks about between the between any two candidates they should have the same result because they are similar. […] I felt like if we do one of these things another metric will also come true”.
A different way to fashion a Custom Fairness metric can be seen in “Equal Opportunity & Counterfactual Fairness”, where different aspects of metrics are combined. Participants who chose this combination aimed to Remove the influence of irrelevant features on AI outcomes through the individual-level Counterfactual Fairness while also trying to inject group-level comparisons with Equal Opportunity by Focusing only on qualified individuals, thus ensuring fair treatment for individuals with a “Good Credit” ground-truth label. As P20 highlighted, “So for ‘Dependents’, I would like a situation of counterfactual fairness, because I think for each individual, like when their dependents [number] change, the AI model should always give them the same prediction, and then as well as that [equal opportunity], for the people that can actually have good credit, that can actually afford it [can be treated fairly]”. In contrast to the previous instance where two metrics are simply assumed to be optimized at the same time, this would result in a new metric altogether, where desirable aspects of the two metrics are integrated.
@@ 4.2.4. Diverse but Strict Thresholds
An important part of fairness decision-making is how much unfairness is acceptable, which we measured using fairness thresholds ranging from 0% to 100%, with higher thresholds indicating greater tolerance for unfairness. Looking at the top Feature-Metric pairings we just described, we found that participants’ fairness demands were notably stringent for existing metrics, while still allowing for Some room for unfairness (13 participants). For example, thresholds were set at on average 6.00% for (Counterfactual Fairness—’Age’) and 4.75% for (Counterfactual Fairness—’Telephone’) while this is even lower for some feature-metric pairings, such as 3.20% for (Equalized Odds—’Foreign Worker’), and 3.50% for (Conditional Statistical Parity—’Dependents’). For Custom Fairness, there are no easy ways to compare between threshold levels but we found that they fluctuated a lot, for example from 0.5% for (“Counterfactual Fairness & Equal Opportunity”—’Gender’) to 40% for (“Outcome Test & Counterfactual Fairness”—’Purpose’) but these might be due to individual differences rather than a consistent pattern. However, the thresholds specified by our participants usually exceeded the legal benchmark of 20% and the common AI development standard of around 10%, regardless of whether the features are protected or non-protected.
Findings summary: In feature-metric pairings, participants tended to choose different metrics for features rather than relying on a “one-size-fits-all” metric selection approach. They also showed a strong interest in creating more complex custom metrics, either to ensure that two or more metrics would be met at the same time or by trying to integrate aspects of different metrics that they liked. Participants seemed to adjust their fairness thresholds based on specific feature-metric pairs, demonstrating stricter fairness demands compared to current practice by AI experts.
@@ 4.3. User Experience: Challenges and Benefits
Our participants reported two main challenges in the process of fairness assessment. First, participants (9) noted that understanding different fairness metrics and terms required considerable time and effort. As one noted, “It took me quite a while to understand the different types of fairness… it also felt that both individual and group fairness should be taken into account… It was also hard to remember the different types and compare at a glance” (P27). Second, participants (3) pointed out that information overload in the prototype and “sometimes difficult to concentrate on where to focus” (P19). As P26 noted, “We have many parameters and they are complexly interrelated. To make a decision, I need to understand them fully”. While participants noted challenges, they all successfully understood and independently completed all three tasks without researcher intervention.
@@
We also collected participants’ System Usability Scale (SUS) (Brooke, 1995) scores to quantitatively assess the usability of our current system (Appendix A.2), thereby informing targeted directions for future design improvements. The overall average SUS score was 60.58, with nearly 65% of participants rating the system between 60 and 100, suggesting moderate usability relative to the desirable benchmark of 68(Brooke, 1995). A closer look at item-level responses provides a more nuanced picture. The SUS includes both “positive” (odd) items and “negative” (even) items: for positive items (0–5 scale, higher is better), participants reported strong agreement that the system is easy to use (24/30), that information is well integrated (30/30), that they felt confident when using it (25/30), and that it is something they would like to use frequently for AI fairness assessment (28/30). Notably, all these items received scores of at least 3, reflecting solid usability signals.
For negative items (0–5 scale, lower is better), very few participants perceived the system as complex (5/30), inconsistent (1/30), or cumbersome (3/30), and these scores fell between 1 and 3, indicating that negative perceptions were limited. While lower scores on questions about needing support contributed to reducing the overall SUS score, this was expected since the system was designed to be used with researcher assistance during the study. Taken together, these results suggest that, despite the inherent complexity of fairness assessment tasks, participants generally found the prototype system straightforward, confidence-enhancing, and well-integrated, offering a strong foundation for further refinement.
@@ 5. Discussion
We acknowledge our study’s limitations and outline future directions. We identify patterns in stakeholders’ fairness decisions and their potential for generalization, propose tool design implications to support these patterns, and call for scaling up stakeholder involvement with appropriate complexity management.
5.1. Limitations
@@
Our qualitative study with 30 stakeholders without AI expertise uncovered rich and detailed insights into their behaviors and underlying reasoning when assessing AI fairness. As an initial step in this direction, we acknowledge several limitations that also point to opportunities for future work. First, while focusing on a single stakeholder type allowed for more targeted analysis, the complexity of behavior is likely to increase when involving a broader range of stakeholder roles (e.g., policymakers, domain experts), backgrounds, and characteristics. Inspired by prior research on fairness perceptions (Nakao et al., 2023), participants’ fairness assessment behaviors may also be shaped by their cultural and societal norms. Future research could expand our methodology with larger and more diverse participant samples, enabling an examination of how fairness assessment varies across roles and cultural contexts.
Second, our study was intentionally situated in the credit rating context. While certain observations, such as specific feature selections, metric choices, and threshold settings, may reflect domain-specific priorities, the underlying patterns of complexity and value-driven reasoning are likely relevant to other AI applications: determining whom to protect (including both legally protected and unprotected features), how to protect (tailoring metrics to specific features), and to what extent (setting strict fairness thresholds). Future research could extend this fairness assessment to diverse domains (e.g., healthcare, hiring, criminal justice) to surface context-specific nuances, as well as to examine how decision-making patterns vary across scenarios with different levels of risk (Haider et al., 2024).
@@
Third, although stakeholders expressed clear preferences for customized fairness metrics, the current prototype system provides limited insight into how such metrics could be operationalized in practice. Our system was designed for outcome fairness and did not support different fairness perspectives, such as affordability or fairness through unawareness. Future research should investigate how stakeholders combine metrics, and explore tools and processes that not only support the creation, validation, and integration of custom metrics into fairness assessment and mitigation frameworks, but also enable fairness assessments that capture diverse perspectives beyond outcome fairness.
5.2. Patterns in AI Fairness Decision-making
In examining how stakeholders assess AI fairness through feature selection and ranking, metric selection, and threshold setting, our scenario-specific data serve as contextual reference. Interestingly, recent research suggested that people’s fairness preferences do not vary significantly across perceived risk levels of different scenarios (Haider et al., 2024). We therefore hypothesize that, although our findings were observed in the credit rating scenario, they are likely to also arise in other high-stakes domains. We identified six recurring decision-making patterns among participants. We compare these six patterns with insights from prior research and analyze their potential generalizability to other high-stakes domains.
@@ 5.2. Patterns in AI Fairness Decision-making
Broader and dynamic feature choices: Stakeholders identify fairness-critical features based on contextual relevance rather than legal protection status, a principle likely to apply in other high-stakes AI domains. Although stakeholders prioritized protected features such as ‘Gender’, which resonates with prior user study (Saxena et al., 2019; Hort et al., 2024) and reflects alignment with established legal and ethical frameworks in the EU (European Union, 2012), the US (Barocas and Selbst, 2016), and the UK (UK Government, 2010), they also elevated non-protected yet contextually less relevant (e.g., to financial ability in our credit rating scenario) features, such as ‘Telephone’, or ‘Purpose’, to high fairness relevance. Our findings demonstrated previous hypotheses that non-protected features can carry fairness weight (Luo et al., 2025; Wachter and Mittelstadt, 2019; Binns, 2020) in high-stakes scenarios.
Preference for individual-by-individual fairness over aggregate group parity: Our results challenge the dominance of group-level metrics in current AI governance (Srivastava et al., 2019; Mehrabi et al., 2021; Binns, 2020). Stakeholders gravitate toward Counterfactual Fairness and other individual-centered metrics. These results are consistent with prior research in credit rating or loan scenarios (Saxena et al., 2019; Nakao et al., 2023; Luo et al., 2025). We believe that this pattern is especially pronounced in contexts where individual circumstances carry significant ethical weight, such as criminal sentencing.

### v2

@@ 4.3.2. Context Matters: Stakeholders Switch Metrics Across Features
Next, we analyzed each participant’s metric selections. We found that 20 participants often switched metrics for different features. For example, 10 participants used 3 different metrics, 8 used 2 metrics, and 2 used 4 metrics. An illustrative example is P7, who ranked and evaluated four distinct features, each aligned with a unique metric: (Consistency — ‘Credit History’); (Custom Fairness— ‘Purpose’); (Equal Opportunity — ‘Foreign Worker’); (Conditional Statistical Parity — ‘Gender’). We observed that participants seemed to first consider their preferred fairness category, i.e., individual or/and group fairness, and then decided on specific metrics within these categories. For example, P16 explained, “so to me the group fairness doesn’t take that much priority because the individual fairness is what matters” which then led to pairing different metrics within the individual fairness category with the selected features. Only six participants followed a “one-size-fits-all” metric selection approach, similar to the practice often adopted by AI experts across both credit scenarios and other AI applications scenarios (de Castro Vieira et al., 2025; Hort et al., 2024), where they chose the same metric (or metric combinations under Custom Fairness) across all of the features they had selected.
@@ 4.3.3. Metrics for Features, and the Reasons Behind Them
We found that each of the top ranked features (‘Gender’, ‘Foreign Worker’, ‘Age’, ‘Telephone’, ‘Dependents’, and ‘Purpose’) were paired with multiple metrics: at least five different metrics in the case of ‘Telephone’, ‘Dependents’, and ‘Purpose’, to seven metrics for ‘Gender’. Since participants chose different metrics for different features, we investigated any patterns within feature-metric pairings, concentrating on the six most selected features and their most frequently selected metrics (Figure 3(b)).
@@
Finally, we found that when predefined metrics did not meet participants’ perceptions, they were trying to define their own fairness metrics. Custom Fairness emerged as one of the most selected metrics for ‘Gender’ (7 participants), ‘Dependents’ (3 participants), and ‘Purpose’ (3 participants). Participants for the most part tried to combine different existing metrics to achieve a more comprehensive fairness assessment, for example by combining “Consistency and Counterfactual Fairness” for ‘Gender’ (2 participants) or “Equal Opportunity and Counterfactual Fairness” for ‘Gender’ (2 participants), ‘Dependents’ (2 participants) and ‘Purpose’ (2 participants).
In constructing these custom metrics, such as “Consistency & Counterfactual Fairness”, participants wanted to meet both individual fairness metrics simultaneously. This, in their mind, would balance Comparison with similar individuals and Remove the influence of irrelevant features. Participants also mentioned the similarity between these two metrics. As P16 explained: “I think it’s interesting because I feel like the two things are connected so I felt like if counterfactual fairness talks about having the same result even when those specific features changed, and then consistency talks about between the between any two candidates they should have the same result because they are similar. […] I felt like if we do one of these things another metric will also come true”.
A different way to fashion a Custom Fairness metric can be seen in “Equal Opportunity & Counterfactual Fairness”, where different aspects of metrics are combined. Participants who chose this combination aimed to Remove the influence of irrelevant features on AI outcomes through the individual-level Counterfactual Fairness while also trying to inject group-level comparisons with Equal Opportunity by Focusing only on qualified individuals, thus ensuring fair treatment for individuals with a “Good Credit” ground-truth label. As P20 highlighted, “So for ‘Dependents’, I would like a situation of counterfactual fairness, because I think for each individual, like when their dependents [number] change, the AI model should always give them the same prediction, and then as well as that [equal opportunity], for the people that can actually have good credit, that can actually afford it [can be treated fairly]”. In contrast to the previous instance where two metrics within the same fairness category are simply assumed to be optimized at the same time, this cross-category combination functions more like a new fairness metric, integrating desirable elements from each metric and potentially posing greater challenges during mitigation.
@@ 4.3.4. Diverse but Strict Thresholds
An important part of fairness decision-making is how much unfairness is acceptable, which we measured using fairness thresholds. Recall that we defined fairness thresholds as the maximum allowable deviation from the optimal fairness outcome of the current metric. Participants set thresholds ranging from 0% to 100%, with higher thresholds indicating greater tolerance for unfairness. Looking at the top Feature-Metric pairings we just described, we found that participants’ fairness demands were notably stringent for existing metrics, while still allowing for Some room for unfairness (13 participants). For example, thresholds were set at on average 6.00% for (Counterfactual Fairness—’Age’) and 4.75% for (Counterfactual Fairness—’Telephone’) while this is even lower for some feature-metric pairings, such as 3.20% for (Equalized Odds—’Foreign Worker’), and 3.50% for (Conditional Statistical Parity—’Dependents’). For Custom Fairness, there are no easy ways to compare between threshold levels but we found that they fluctuated a lot, for example from 0.5% for (“Counterfactual Fairness & Equal Opportunity”—’Gender’) to 40% for (“Outcome Test & Counterfactual Fairness”—’Purpose’) but these might be due to individual differences rather than a consistent pattern. However, the thresholds specified by our participants usually exceeded thresholds often used by AI experts, such as the legal benchmark of 20% deviation (Feldman et al., 2015) or a common AI development standard of around 10% deviation (Gupta et al., 2021), regardless of whether the features are protected or non-protected.
Findings summary: In feature-metric pairings, participants tended to choose different metrics for features rather than relying on a “one-size-fits-all” metric selection approach. They also showed a strong interest in creating more complex custom metrics, either to ensure that two or more metrics would be met at the same time or by trying to integrate aspects of different metrics that they liked. Participants seemed to adjust their fairness thresholds based on specific feature-metric pairs, demonstrating high expectations for fairness and low tolerance for unfairness.
@@ 4.4. User Experience: Challenges and Benefits
Our participants reported two main challenges in the process of fairness assessment. First, 9 participants noted that understanding different fairness metrics and terms required considerable time and effort. As one noted, “It took me quite a while to understand the different types of fairness… it also felt that both individual and group fairness should be taken into account… It was also hard to remember the different types and compare at a glance” (P27). This challenge was not unexpected: participants had no AI expertise, yet assessing AI fairness required them to understand each metric and decide which one best fit their fairness assessment needs. Second, only 3 participants pointed out that information overload in the prototype made it “sometimes difficult to concentrate on where to focus” (P19). As P26 noted, “We have many parameters and they are complexly interrelated. To make a decision, I need to understand them fully”.
These challenges highlight a tool design tension: providing enough information for stakeholders to understand and assess AI fairness without causing information or cognitive overload. To tackle these challenges, we observed that for metric exploration, participants typically moved through the interface from top to bottom: reading the textual definition, checking the visualization and calculation, and then understanding instance-level explanations. They deepened their understanding by selecting the explanation formats that suited them and lingering on the explanations to think them through. When deciding on a metric, participants often switched across UI pages to compare the pros and cons of different metrics, align the options with their needs, and consider whether to create custom fairness metrics. We will outline potential tool improvements in Section 5.3.
While participants noted challenges, they all successfully understood and independently completed all three tasks without researcher intervention in their fairness decision-making. Because participants thought aloud and the researcher asked follow-up questions when needed, their explanations made clear how they interpreted the scenario, understood different features and fairness metrics, compared different fairness metrics, and articulated why particular metrics aligned with their concerns for specific features. These real-time observations enabled us to verify participants’ correct understanding, while our thematic analysis captures how they reasoned with this understanding to make value-driven fairness assessment choices.
@@
We also collected participants’ System Usability Scale (SUS) (Brooke, 1995) scores to quantitatively assess the usability of our current system (Appendix A.3), thereby informing targeted directions for future design improvements. The overall average SUS score was 61.83, with nearly 75% of participants rating the system between 60 and 100, suggesting moderate usability relative to the desirable benchmark of 68 (Brooke, 1995). A closer look at item-level responses provides a more nuanced picture. The SUS includes both “positive” (odd) items and “negative” (even) items: for positive items (0–5 scale, higher is better), participants reported strong agreement that the system is easy to use (23/26), that information is well integrated (26/26), that they felt confident when using it (23/26), and that it is something they would like to use frequently for AI fairness assessment (24/26). Notably, all these items received scores of at least 3, reflecting solid usability signals.
For negative items (0–5 scale, lower is better), very few participants perceived the system as complex (4/26), inconsistent (1/26), or cumbersome (2/26), and these scores fell between 1 and 3, indicating that negative perceptions were limited. While lower scores on questions about needing support contributed to reducing the overall SUS score, this was expected since the system was designed to be used with researcher assistance for technical system usage, available when needed during the study. Taken together, these results suggest that, despite the inherent complexity of fairness assessment tasks, participants generally found the prototype system straightforward, confidence-enhancing, and well-integrated, offering a strong foundation for further refinement.
@@ 5. Discussion
We acknowledge our study’s limitations and outline future directions. We identify patterns in stakeholders’ fairness assessment behavior and propose tool design implications to support stakeholders in assessing AI outcome fairness. We also provide possible approaches to manage the complexities of stakeholders’ fairness decisions.
5.1. Limitations and Future Work
Our qualitative study with 26 stakeholders without AI expertise uncovered rich and detailed insights into their behaviors and underlying reasoning when assessing AI outcome fairness. As an initial step in this direction, we acknowledge several limitations that also point to opportunities for future work.
@@ 5.1. Limitations and Future Work
First, while focusing on a single stakeholder type allowed for more targeted analysis, the complexity of behavior is likely to increase when involving a broader range of stakeholder roles (e.g., policymakers, domain experts), backgrounds, and characteristics. Inspired by prior research on fairness perceptions (Nakao et al., 2023), participants’ fairness assessment behaviors may also be shaped by their cultural and societal norms. Future research could expand our methodology with larger and more diverse participant samples, enabling an examination of how fairness assessment varies across roles and cultural contexts. Moreover, we observed the emergence of differences between lay stakeholders’ (decision subjects) fairness assessment decision-making and the prevailing practices of AI experts as reflected in prior work (de Castro Vieira et al., 2025; Hort et al., 2024; Verma and Rubin, 2018; Rabonato and Berton, 2025). Future work could involve both AI experts and lay stakeholders to collaboratively explore ways to reconcile differences and build consensus toward more inclusive AI fairness.
Second, our study was intentionally situated in the credit rating context, and although the German Credit dataset is real-world, it is relatively dated compared with more modern datasets that contain richer features, missing values, or multi-modal data. This might constrain the breadth of fairness assessment behaviors we were able to observe. Nonetheless, the underlying patterns of complexity and value-driven reasoning we uncovered in this scenario may still advance our understanding of stakeholders’ needs when assessing fairness and offer useful insights for HCI-driven AI fairness design. Future research could extend this fairness assessment to diverse domains (e.g., healthcare, hiring, criminal justice) to surface context-specific nuances, as well as examine how decision-making patterns vary across scenarios with different levels of risk (Haider et al., 2024).
@@
Third, although stakeholders expressed clear preferences for customized fairness metrics, the current prototype system does not support direct custom fairness input in an operational formulation, thereby providing limited insight into operationalization and mitigation in practice. Future research should investigate tools and processes that not only help stakeholders articulate custom fairness definitions that reflect their expectations, but also translate these lay concepts into operationalizable metric formulations, thereby supporting the validation and integration of custom metrics into fairness assessment and mitigation workflows.
Additionally, although we found other fairness perspectives held by our participants, our prototype was designed specifically for outcome fairness and did not support fairness assessments grounded in other perspectives, such as affordability fairness or fairness through unawareness, which required specifying which features to include or exclude during model development and evaluating fairness on that basis. Future work should investigate how stakeholders could be supported in articulating different fairness perspectives and examine what factors might shape these perspective differences, such as demographic factors.
5.2. Patterns in Stakeholders’ AI Fairness Assessment Decision-making
We supported participants in assessing AI outcome fairness through feature selection (access to both protected and unprotected features), ranking (to express fairness priorities), metric selection (with group, individual, and custom fairness options), and threshold setting (at the feature–metric level). As fairness is highly contextual (Kim et al., 2025) and changes in domain or even dataset can lead to different selection outcomes, we therefore do not focus on participants’ specific choices. Instead, we carefully identify six recurring decision-making patterns that capture overall preferences in how participants assessed fairness across features, metrics, thresholds, and priorities. We next discuss these six patterns alongside prior research and reflect on how they may offer some tentative insights into how lay stakeholders assess AI outcome fairness.
5.2.1. Broad feature selection with expanded fairness considerations
@@
Although participants prioritized protected features such as ‘Gender’, which resonates with prior user studies (Saxena et al., 2019; Hort et al., 2024) and reflects alignment with established legal and ethical frameworks in the EU (European Union, 2012), the US (Barocas and Selbst, 2016), and the UK (UK Government, 2010), they also elevated non-protected yet contextually less relevant (e.g., to financial ability in our credit rating scenario) features, such as ‘Telephone’, or ‘Purpose’, to high fairness relevance. These observations lend partial support to previous hypotheses that non-protected features might carry fairness weight in high-stakes scenarios (Luo et al., 2025; Wachter and Mittelstadt, 2019; Binns, 2020). Taken together, our findings imply that when stakeholders identify fairness-related features, they are likely to consider contextual relevance and the potential bias a feature may introduce, rather than relying on legally protected characteristics. By contrast, AI experts’ current practice in credit contexts (de Castro Vieira et al., 2025) and other AI domains (Lalor et al., 2024) often relies on a limited set of legally protected features. The broader feature considerations shown by stakeholders suggest that this area may warrant more attention.
5.2.2. Preference for fine-grained fairness assessment
From participants’ overall metric choices, we observed a preference toward fine-grained, comprehensive fairness measurement, in contrast to coarse group-based parity. This preference was reflected, in part, through participants’ frequent selection of Counterfactual Fairness, which they described as a way to treat people fairly at the individual level and to remove the influence of irrelevant features when AI makes decisions. In contrast, coarse group-based measures were relatively unpopular. For example, no participant selected Demographic Parity. When participants pursued group fairness for features, they leaned toward more fine-grained forms of group assessment. For example, more people preferred metrics like Conditional Statistical Parity for customizing comparison groups and making feature-specific, refined group evaluation, compared to Demographic Parity or Equalized Odds. This pattern resonates with similar metric preferences in prior studies involving lay stakeholders in credit rating, loan decision-making, and child maltreatment risk prediction, which also showed that stakeholders examined fairness at a more fine-grained level rather than relying solely on coarse group statistics (Saxena et al., 2019; Nakao et al., 2023; Luo et al., 2025; Cheng et al., 2021). Recent work in personnel selection also reports a similar preference for ’concrete’ fairness: stakeholders often felt the need to supplement group-based fairness metrics with more individual-level examinations, noting that coarse group statistics could overlook important nuances (Müller et al., 2025). In contrast, current practice among AI experts tends to simplify fairness assessment to group-statistical comparisons, such as Demographic Parity, because they offer greater operational practicality, align with legal requirements, and facilitate comparability, especially during bias mitigation (de Castro Vieira et al., 2025; Hort et al., 2024; Verma and Rubin, 2018; Mehrabi et al., 2021; Binns, 2020).
5.2.3. Favoring custom metrics
Given that people may hold different expectations of fairness shaped by their lived experiences and backgrounds (Landers and Behrend, 2023; Lee et al., 2020), we provided a larger list of existing metrics for participants to choose from (8 metrics covering group and individual fairness), compared to most user studies (Haider et al., 2024; Cheng et al., 2021; Saxena et al., 2019). Yet, participants still expressed a clear interest in designing custom metrics tailored to their needs. Our results show that participants often combined multiple existing metrics to construct their own fairness notions, which also reflects the need for more fine-grained and comprehensive fairness assessments discussed above. This an area that has received relatively little attention so far. While previous work has speculated on the need for expressing these custom fairness metrics (Nakao et al., 2023), our work is the first evidence that this is required by stakeholders. This finding may help guide future work by supporting stakeholders to express custom fairness metrics, beyond simply choosing one metric from a set of predefined options.
At the same time, there is a need for AI experts to explore ways to handle combinations of fairness metrics, as no unified guidance exists for metric selection, and a single metric rarely captures the complexity of real-world fairness needs (Mehrabi et al., 2021). From this perspective, a human-centered approach to fairness may offer a promising direction.
However, such metric combinations create operational challenges for AI experts, as many fairness metrics are mathematically incompatible and optimizing one often degrades another (Mehrabi et al., 2021; Binns, 2020; Dwork et al., 2012; Pessach and Shmueli, 2022). In practice, such custom “combinations” of fairness metrics could be implemented in several ways: through a logical operator ‘AND’ (all selected metrics must be satisfied), a weighted multi-objective approach (metrics contribute with assigned weights) (Zhang et al., 2020), or a constraint hierarchy (higher-priority metrics take precedence) (Luo et al., 2025). When tensions arise among the selected metrics, the trade-off could be visualized, such as through a Pareto frontier, helping AI experts or stakeholders navigate the space of workable compromises in deployment (Yu et al., 2020).
5.2.4. Tailoring metrics to features